The following information is available to determine the budget requirements of Medico Limited for
January 2021:
1. Projected sales of Product Med for January and February 2021:
January February
15 000 units at R30 each 12 000 units at R32 each

2. The policy of Medico Limited is to maintain a closing inventory equal to 40% of the budgeted sales
of the following month.
3. Two (2) labour hours are required to produce one unit of Product Med and the rate per hour is
R10.
4. Overheads are projected at a rate of R3 per direct labour hour.
5. The estimated total production costs per unit (including the costs above) are R40 for January 2021
and R38 for December 2020.
REQUIRED
Use the information provided above to prepare the following budgets for January 2021:
1.1 Sales Budget (units and Rands)
1.2 Production Budget (units)
1.3 Direct Labour Budget (Rands)
1.4 Overheads Budget (Rands)
1.5 Cost of Sales Budget(Rands)
1
Expert's answer
2020-05-12T17:21:18-0400

1.1 "15000\times30=450 000"

1.2 "15000\times0.4=6000" inventory balance at the beginning of January

15000 - sales, "12000*0.4=4800 -" inventory balance at the end of January

hence the volume of production for January :6000+x-15000=4800, x=13800

1.3 "13 800\times2\times10=276 000"

1.4 "13800\times2\times3=82 800"

1.5 "40\times 13800+6000\times38-40\times4800)=552 000+228 000-192 000=588 000"